# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v2

@@ 5.2.1 Implementation of ODEBlock
Fig. 7 illustrates the ODEBlock implementation. Since the residual connection of ResBlocks is replaced with ODEBlock, the shortcut connection is not included in the ODEBlock. In the ODEBlock, a numerical computation to solve the ordinary differential equation in Eq. (17) is carried out as described in Sec. 3.2. As shown in Table 1, ODEBlocks consist of several modules, such as Add time, DSC, BatchNorm, and ReLU. The weight buffer size for each ODEBlock is determined by DSC and BatchNorm (Add time and ReLU do not have parameters). DSC consists of point-wise and depth-wise convolution layers; thus, the weight buffer size for each DSC is $Ch_{in}\times K\times K+Ch_{out}\times Ch_{in}$. BatchNorm requires two parameters precomputed by software for each channel; thus, the weight buffer size for each BatchNorm is $2\times Ch$. Assuming model sizes in Table 1, the weight buffer sizes for ODEBlock1 and ODEBlock2 are 9,618 and 35,858, respectively.
As described in Sec. 3.2, the number of iterations in ODEBlock is given as a hyperparameter so that we can determine the time interval for the integration. The number of iterations for each ODEBlock is specified through respective control registers. ODEBlock keeps track of a time variable $t$ ($t_{j}=t_{0}+jh$ in Eq. (17)) during iterations ($j=0,\ldots,C-1$). The module “Add time” in Table 1 is to insert an additional channel filled with a time variable $t$ to the layer input; thus, this increases the number of output channels by one. In addition to on-chip buffers for weight parameters, extra buffers are allocated to retain the input to the block for skip connection and the output of each layer.
@@ 5.2.2 Implementation of DSBlock
Fig. 8 illustrates the DSBlock implementation. The main purpose of DSBlock is to reduce the feature map size by half in the model. Thus, DSBlocks are not iterated unlike ODEBlocks, and Neural ODE is not applied to the DSBlock in our implementation. DSBlocks consist of several modules, such as convolution, BatchNorm, and ReLU. The weight buffer size for each DSBlock is determined by the convolution and BatchNorm. The weight buffer size for each convolution layer is $Ch_{out}\times Ch_{in}\times K\times K$. The weight buffer size for each BatchNorm is $2\times Ch$ as well as ODEBlock. Assuming model sizes in Table 1, the weight buffer sizes for DSBlock1 and DSBlock2 are 237,047 and 925,943, respectively.
@@ 5.2.3 Implementation of MHSABlock
Fig. 9 illustrates the MHSABlock implementation. As well as the ODEBlock, the shortcut connection is not included in the MHSABlock. In the MHSABlock, a spatial convolution layer of the original ResNet building block is replaced with MHSA. MHSA (Fig. 4) is executed in the following sequence: (1) Three weights $\mathbf{W}_{k}$, $\mathbf{W}_{q}$, and $\mathbf{W}_{v}$, along with pre-trained $\mathbf{R}_{w}$ and $\mathbf{R}_{h}$ for positional encoding, are transferred to the on-chip buffers. Then, (2) $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{QK}^{\top}$ are calculated sequentially for input $\mathbf{X}$. (3) For positional encoding, $\mathbf{R}=\mathbf{1}_{w}^{\top}\mathbf{R}_{h}+\mathbf{1}_{h}\mathbf{R}_{w}$ and $\mathbf{QR}^{\top}$ are calculated sequentially. (4) $\mathbf{QK}^{\top}$ is added to the positional encoding $\mathbf{QR}^{\top}$ (i.e., $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$). Afterward, (5) $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$ is divided by a scaling factor $\sqrt{D_{h}}$. (6) ReLU is applied for $\frac{\mathbf{Q}\mathbf{K}^{\top}+\mathbf{Q}\mathbf{R}^{\top}}{\sqrt{D}_{h}}$ to acquire an attention $\mathbf{A}$. Subsequently, (7) $\mathbf{V}$ is calculated for input $\mathbf{X}$. (8) A dot product of $\mathbf{A}$ and $\mathbf{V}$ is calculated. Finally, (9) LayerNorm is applied for the output of MHSA.
The MHSABlock consists of multiple modules, such as point-wise convolution, BatchNorm, and MHSA. The MHSA has weight parameters $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. In addition, those for the positional encoding and LayerNorm are needed. The weight buffer size for $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ is $3\times Ch\times Ch$ in total. That for the positional encoding is $Ch\times H\times W$. LayerNorm also requires two parameters precomputed by software for each feature; thus, the weight buffer size for the LayerNorm is $2\times Ch\times H\times W$. Assuming model sizes in Table 1, the weight buffer size for MHSABlock is 64,433.
@@
5.3 Quantization of layers
$32\times(2^{n}\times 9+2)<(32-n)\times N_{p}$ (30)
$\frac{32\times N_{p}}{n\times N_{p}+n\times 2^{n}\times 9+32\times 2}.$ (31)
In our FPGA implementation, the quantization is selectively applied to some specific blocks by considering the accuracy and parameter size. In Sec. 6.2.3, ablation studies to investigate the quantization will be conducted and show that the model where the quantization is applied to DSBlocks and MHSABlock (denoted as “Block3” in Table 5) strikes a good balance between the accuracy and parameter size. Based on this result, 4-bit or 8-bit quantization is applied to DSBlocks and MHSABlock to implement dedicated IP cores.
5.4 Degree of parallelism in each layer
In our FPGA implementation, all the weight parameters processed with the PL part are stored in on-chip buffers; thus, the main bottleneck in execution time lies in the computation rather than the data transfer. Thus, parallelization is applied to all the layers in Table 1 to speed up the computation as much as the FPGA resources such as BRAM and DSP allow. Based on simulation results with Vitis HLS, a higher degree of parallelism is applied to layers with longer execution time. Table 2 shows the degree of parallelism and the number of cycles for each layer. In this table, SC, DW, PW, and BNReLU stand for shortcut connection, depth-wise, point-wise, and BatchNorm and ReLU layers, respectively. Note that the same degree of parallelism is applied for both the 4-bit and 8-bit quantized implementations.
@@ 6.2.1 Parameter size and computational cost
Table 3 shows the parameter size and GFLOPs (giga floating-point operations) of the proposed and counterpart models with and without MHSA mechanism. As shown, the proposed model uses the smallest number of parameters among these lightweight models, thanks to the benefits of Neural ODE. BoTNet50 has 1.24$\times$ less parameters than ResNet50, because convolutions in the last three ResBlocks are replaced by MHSA. The proposed model achieves 14.96$\times$ reduction from BoTNet50 by combining MHSA and Neural ODE; as described in Sec. 3, the use of Neural ODE allows to compress $C$ ResBlocks into one ODEBlock, yielding $C$-fold reduction. In this case, $C$, i.e., the number of ODE solver iterations, is set to 10. By this model compression, the proposed model requires a high computational demand relative to parameter size. From the table, it can be observed that GFLOPs per million parameters in ODENet and the proposed model (0.17–0.48) are mostly larger compared to the counterpart models (0.0039–0.20). These models are thus well-suited for hardware accelerators in edge devices with limited memory capacity.
@@ 6.2.2 Classification accuracy
Learning rate Model 8-bit acc (%) 4-bit acc (%)
@@
(a) [5e-3,1e-5] DS only 81.29 80.06
Block1 76.13 73.21
Block2 75.81 76.51
Block3 81.56 80.78
Block1+2 62.40 65.01
Block1+3 71.76 75.20
Block2+3 77.11 76.21
(b) [5e-4,1e-5] DS only 81.46 80.38
Block1 77.61 78.35
Block2 77.61 76.31
Block3 80.40 79.83
Block1+2 72.30 64.71
Block1+3 76.30 77.15
Block2+3 77.38 78.30
(c) [1e-4,1e-5] DS only 80.97 80.23
Block1 76.59 76.98
Block2 75.94 76.45
Block3 79.10 80.53
Block1+2 66.19 65.6
Block1+3 73.70 75.24
Block2+3 72.26 76.10
@@ 6.2.3 Effects of quantization
To reduce the memory footprint and implement the proposed model on modest-sized FPGAs, we want to apply quantization to as many layers as possible. On the other hand, since there is a trade-off between the degree of quantization and accuracy, there is a risk of significant accuracy degradation if quantization is applied to all layers without careful consideration. Here, we conduct ablation studies to investigate the impact of learning rates on the quantized blocks. Different combinations of quantized blocks are evaluated. In this experiment, 4-bit and 8-bit quantizations are applied to both weights and activations for all the target blocks. It should be noted that quantization is always applied to DSBlocks that are used only once in a single inference. Quantization on these blocks has less impact on accuracy. DSC is applied to ODEBlocks only. The initial and minimum learning rates for the cosine annealing scheduler are set to [5e-3, 1e-5], [5e-4, 1e-5], and [1e-4, 1e-5].
Table 5 shows the accuracy of each combination of quantized blocks. The accuracy before the quantization is also shown. In this table, “DS only” indicates that the quantization is applied to DSBlocks only. “Block1”, “Block2”, and “Block3” refer to the proposed models with the quantized ODEBlock1, ODEBlock2, and MHSABlock, respectively. “Block1+2” refers to the model with both ODEBlock1 and ODEBlock2 quantized. The best accuracy is achieved when the quantization is applied to DSBlocks only and the learning rate settings is [5e-3, 1e-5] for both 4-bit and 8-bit implementations. Considering that Block3 gives almost the same performance as the DS only setting and achieves higher accuracy than Block1 and Block2, MHSABlock is more tolerant to quantization than ODEBlocks. In addition, while ODEBlock2 contains more parameters than ODEBlock1, Block2 obtains better results than Block1. This implies that the having more parameters in a block does not necessarily increase its sensitivity to quantization.
@@ 6.2.4 FPGA resource utilization
Table 6 presents the FPGA resource utilization of the proposed model with 4-bit and 8-bit quantization applied to all building blocks. Thanks to the quantization and the 14.96$\times$ parameter size reduction by Neural ODE, the feature extraction part of the proposed model fits within the on-chip BRAM/URAM. The 4-bit or 8-bit quantization is applied to both weights and activations for all the target blocks. In the 4-bit and 8-bit implementations, BRAMs are preferentially consumed for weights and activations, and URAMs are utilized for the portions that exceed the BRAM capacity. Thus, BRAM utilizations of the 4-bit and 8-bit implementations are almost the same while their URAM utilizations differ. As a result, compared to the 8-bit quantization, the URAM utilization is further reduced by 17.7% with the aggressive 4-bit quantization. In the case of the Zynq UltraScale+ MPSoC ZCU104 platform, the 8-bit FPGA implementation already meets the resource constraints, and the 8-bit implementation is better than the 4-bit implementation in terms of accuracy. The number of DSP slices is the same in the 4-bit and 8-bit implementations. That is, a single DSP slice is used for a single operation regardless of the quantized bit widths. The benefit of the 4-bit implementation over the 8-bit one is the parameter size reduction.
@@ 6.2.5 FPGA inference accuracy
Here, our 4-bit and 8-bit FPGA implementations are executed on the ZCU104 platform and evaluated in terms of inference accuracy. As shown in Table 5, our simulation results show that the Block3 model where the quantization is applied to DSBlocks and MHSABlock with learning rate of [5e-3, 1e-5] strikes a good balance between the accuracy and parameter size. As a result, 4-bit and 8-bit versions of the Block3 model are implemented on the ZCU104 platform as mentioned in Sec. 5.3. Please note that float32 is used for blocks other than DSBlocks and MHSABlock in the software simulation while a fixed-point number format is used for the other blocks in the FPGA implementation. To see this subtle difference, the simulation and on-board results of the 4-bit and 8-bit implementations are compared in terms of inference accuracy. The results in Table 8 demonstrate that the accuracy loss of the FPGA implementation is quite small in both the 4-bit and 8-bit cases.
Table 8: Top-1 accuracy in software and FPGA board
Platform Model 8-bit acc (%) 4-bit acc (%)
Simulation Block3 81.56 80.78
FPGA board Block3 81.53 79.18

### v3

@@ 5.2.1 Implementation of ODEBlock
Fig. 7 illustrates the ODEBlock implementation. Since the residual connection of ResBlocks is replaced with ODEBlock, the shortcut connection is not included in the ODEBlock. In the ODEBlock, the ordinary differential equation in Eq. (17) is solved numerically as described in Sec. 3.2. As shown in Table 1, ODEBlocks consist of several modules, such as Add time, DSC, BatchNorm, and ReLU. The weight buffer size for each ODEBlock is determined by DSC and BatchNorm (Add time and ReLU do not have parameters). DSC consists of point-wise and depth-wise convolution layers; thus, the weight buffer size for each DSC is $Ch_{in}\times K\times K+Ch_{out}\times Ch_{in}$, where $Ch_{in}$ and $Ch_{out}$ represent the numbers of input and output channels of the feature maps, respectively. BatchNorm requires two parameters precomputed by software for each channel; thus, the weight buffer size for each BatchNorm is $2\times Ch$. In our case (Table 1), the weight buffer sizes for ODEBlock1 and ODEBlock2 are 9,618 and 35,858, respectively.
As described in Sec. 3.2, the number of iterations $C$ in ODEBlock is given as a hyperparameter and specified through a control register. The integration step size $h$ is determined as $1/C$. ODEBlock keeps track of a time variable $t$ ($t_{j}=t_{0}+jh$ in Eq. (17)) during iterations ($j=0,\ldots,C-1$). The module “Add time” in Table 1 is to insert an additional channel filled with a time variable $t$ to the layer input; this increases the number of output channels by one. In addition to on-chip buffers for weight parameters, extra buffers are allocated to retain the input to the block for skip connection and the output of each layer.
@@ 5.2.2 Implementation of DSBlock
Fig. 8 illustrates the DSBlock implementation. The main purpose of DSBlock is to reduce the feature map size by half in the model. DSBlocks are not iterated unlike ODEBlocks, and Neural ODE is not applied to the DSBlock in our implementation. DSBlocks consist of several modules, such as convolution, BatchNorm, and ReLU. The weight buffer size for each DSBlock is determined by the convolution and BatchNorm. The weight buffer size for each convolution layer is $Ch_{out}\times Ch_{in}\times K\times K$. The weight buffer size for each BatchNorm is $2\times Ch$. In our case (Table 1), the weight buffer sizes for DSBlock1 and DSBlock2 are 237,047 and 925,943, respectively.
@@ 5.2.3 Implementation of MHSABlock
Fig. 9 illustrates the MHSABlock implementation. As in the ODEBlock, the shortcut connection is not included in the MHSABlock. In the MHSABlock, a spatial convolution layer of the original ResNet building block is replaced with MHSA. In addition to the convolution weights and batch normalization parameters, the three weights ($\mathbf{W}_{k}$, $\mathbf{W}_{q}$, $\mathbf{W}_{v}$) and pre-trained parameters for positional encoding ($\mathbf{R}_{w}$, $\mathbf{R}_{h}$) are stored on-chip. MHSA (Fig. 4) is executed as follows: (1) The query $\mathbf{Q}$, key $\mathbf{K}$, and product $\mathbf{QK}^{\top}$ are calculated from the input $\mathbf{X}$. (2) Then, $\mathbf{QR}^{\top}$ is calculated and added to $\mathbf{QK}^{\top}$, where $\mathbf{R}=\mathbf{1}_{w}^{\top}\mathbf{R}_{h}+\mathbf{1}_{h}\mathbf{R}_{w}$ denotes the positional encoding. (3) The result $\mathbf{QK}^{\top}+\mathbf{QR}^{\top}$ is divided by a scaling factor $\sqrt{D_{h}}$, and ReLU is applied to obtain an attention $\mathbf{A}$ (Eq. 28). (4) The value $\mathbf{V}$ and product $\mathbf{A}\mathbf{V}$ (Eq. 7) are calculated from the input $\mathbf{X}$, and LayerNorm is applied to produce the final output of MHSA (Eq. 29).
The MHSABlock consists of multiple modules, such as Add time, point-wise convolution, BatchNorm, and MHSA. MHSA has three weight parameters ($\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$) as well as those for the positional encoding and LayerNorm. The weight buffer size for $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ is $3\times Ch\times Ch$ in total and that for the positional encoding is $Ch\times H\times W$. LayerNorm also requires two parameters precomputed by software for each feature; the weight buffer size is $2\times Ch\times H\times W$. In our case (Table 1), the weight buffer size for MHSABlock is 64,433.
@@ 65 $\rightarrow$ 256 $1\times 1$ Conv
5.3 Quantization of layers
$32\times(2^{n}\times 9+2)<(32-n)\times N_{p}$ (30)
$\frac{32\times N_{p}}{n\times N_{p}+n\times 2^{n}\times 9+32\times 2}.$ (31)
In our FPGA implementation, the quantization is selectively applied to some specific blocks by considering the accuracy and parameter size. In Sec. 6.2.3, ablation studies are conducted to investigate the effect of quantization and show that the model with DSBlocks and MHSABlock quantized (denoted as “DS+Block3” in Table 5) strikes a good balance between the accuracy and parameter size. Based on this result, 4-bit or 8-bit quantization is applied to DSBlocks and MHSABlock to implement dedicated IP cores.
5.4 Local memory size
The weight buffer size is 1,272,899 (9,618 + 35,858 + 237,047 + 925,943 + 64,433) in total as mentioned in Secs. 5.2.15.2.3. It slightly differs from the number of parameters (1.26M) reported in Table 3 due to the additional parameters for the quantization (see Sec. 5.3). In addition, intermediate buffers are used to store output feature maps, and their size is 261,648 in total. As a result, the total buffer size is 1,534,547. In the DS+Block3 model, the quantized weight parameters (i.e., those in DSBlocks and MHSABlock) are represented as 4-bit or 8-bit integers. The non-quantized parameters are represented as 16-bit fixed-point number format (integer part: 4-bit) in the PL part. The LLT-quantized activations are also represented as 4-bit or 8-bit integers. The other activations are represented as 20-bit fixed-point number format (integer part: 10-bit) in the PL part. As a result, the local memory sizes for the 4-bit and 8-bit quantized FPGA implementations are 1.23MB and 1.88MB, respectively. Since 11Mb Block RAMs and 27Mb Ultra RAMs are available on the ZCU104 platform, these local memories can be placed on-chip.
5.5 Degree of parallelism in each layer
In our FPGA implementation, all the weight parameters used in the PL part are stored on-chip; thus, the performance bottleneck lies in the computation rather than the data transfer. Thus, parallelization is applied to all the layers in Table 1 to speed up the computation while fully utilizing the FPGA resources such as BRAM and DSP. Based on simulation results from Vitis HLS, a higher degree of parallelism is applied to layers with longer execution time. Table 2 shows the degree of parallelism and the number of cycles for each layer. In this table, SC, DW, PW, and BNReLU stand for shortcut connection, depth-wise, point-wise, and BatchNorm and ReLU layers, respectively. Note that the same degree of parallelism is applied for both the 4-bit and 8-bit quantized implementations.
@@ 6.1.2 Model training
6.1.3 On-board execution
The target FPGA board is the Zynq UltraScale+ MPSoC ZCU104. As mentioned in Sec. 5, the PL part is in charge of ODEBlock1 to MHSABlock, while the PS part is in charge of the pre- and post-processing. The PL part is reconfigured as the proposed IP core. Pynq Linux 2.7 is running on the PS part. Python 3.8.2 and PyTorch 1.13.1 are also running on the PS part. The trained model and test dataset are downloaded to the PS part and then used by the successive blocks in the PL part. The execution times of the PS and PL parts are measured by wall-clock time.
@@ 6.2.1 Parameter size and computational cost
Table 3 shows the parameter size and GFLOPs [^3] GFLOPs highly depend on the input image size and those for STL10 (image size: 96$\times$96) reported in Table 3 are smaller than those for ImageNet (image size: 224$\times$224). (giga floating-point operations) of the proposed and counterpart models with and without MHSA mechanism for the STL10 dataset. As shown, the proposed model uses the smallest number of parameters among these lightweight models, thanks to the benefits of Neural ODE. BoTNet50 has 1.24$\times$ less parameters than ResNet50, because convolutions in the last three ResBlocks are replaced by MHSA. The proposed model achieves 14.96$\times$ reduction from BoTNet50 by combining MHSA and Neural ODE; as described in Sec. 3, the use of Neural ODE allows to compress $C$ ResBlocks into one ODEBlock, yielding $C$-fold reduction. In this case, $C$, i.e., the number of ODE solver iterations, is set to 10. By this model compression, the proposed model requires a high computational demand relative to parameter size. From the table, it can be observed that GFLOPs per million parameters in ODENet and the proposed model (0.17–0.48) are mostly larger compared to the counterpart models (0.0039–0.20). These models are thus well-suited for hardware accelerators in edge devices with limited memory capacity.
@@ 6.2.2 Classification accuracy
Learning rate Model 8-bit quant. (%) 4-bit quant. (%)
@@
(a) [5e-3,1e-5] DS only 80.97$\pm$0.08 81.48$\pm$0.21
Block1 76.06$\pm$0.08 76.32$\pm$3.88
Block2 78.18$\pm$0.17 77.49$\pm$0.12
Block3 81.23$\pm$0.07 80.96$\pm$0.01
DS+Block3 81.23$\pm$0.27 78.84$\pm$0.11
(b) [5e-4,1e-5] DS only 80.83$\pm$0.15 81.06$\pm$0.05
Block1 77.51$\pm$0.04 78.48$\pm$0.03
Block2 78.71$\pm$0.03 78.69$\pm$0.00
Block3 80.48$\pm$0.03 80.68$\pm$0.03
DS+Block3 80.48$\pm$0.17 80.03$\pm$0.13
(c) [1e-4,1e-5] DS only 79.70$\pm$0.05 80.27$\pm$0.20
Block1 75.15$\pm$0.04 77.16$\pm$0.02
Block2 77.25$\pm$0.00 77.95$\pm$0.01
Block3 79.61$\pm$0.00 80.15$\pm$0.00
DS+Block3 76.53$\pm$0.20 79.10$\pm$0.18
@@ 6.2.3 Effects of quantization
To reduce the memory footprint and implement the proposed model on modest-sized FPGAs, we want to apply quantization to as many layers as possible. On the other hand, since there is a trade-off between the degree of quantization and accuracy, there is a risk of significant accuracy degradation if quantization is applied to the model without careful consideration. Here, we conduct ablation studies to investigate the impact of learning rates and quantized blocks. Different combinations of the learning rates and quantized blocks are evaluated. In this experiment, 4-bit and 8-bit quantizations are applied to both weights and activations for all the target blocks. The initial and minimum learning rates for the cosine annealing scheduler are set to [5e-3, 1e-5], [5e-4, 1e-5], and [1e-4, 1e-5]. DSC is applied to ODEBlocks only.
Table 5 shows the accuracy of the proposed model when the 4-bit or 8-bit quantization is applied to different blocks. The accuracy without the quantization is also shown. In this table, “DS only” indicates that the quantization is applied to DSBlocks only. “Block1”, “Block2”, and “Block3” refer to the proposed models with the quantized ODEBlock1, ODEBlock2, and MHSABlock, respectively. “DS+Block3” indicates that the quantization is applied to DSBlocks and MHSABlock. Each model is trained with different random seeds and tested three times. Mean accuracies and their standard deviations are shown in the table. These results are also visualized in Figs. 11 and 12. For the 8-bit implementation, the best accuracy is achieved when the quantization is applied to MHSABlock only and the learning rate setting is [5e-3, 1e-5]. For the 4-bit implementation, the best accuracy is achieved when the quantization is applied to DSBlocks only. Considering that DS+Block3 gives almost the best accuracy for the 8-bit implementation while reducing the parameter size compared to DS only and Block3, in this paper we implement the DS+Block3 model on the FPGA, as evaluated below.
@@ 6.2.4 FPGA resource utilization
Table 6 presents the FPGA resource utilization of the DS+Block3 model for the 4-bit and 8-bit implementations. Thanks to the quantization and the 14.96$\times$ parameter size reduction by Neural ODE, the feature extraction part of the proposed model fits within the on-chip BRAM/URAM. The 4-bit or 8-bit quantization is applied to both weights and activations for DSBlocks and MHSABlock. In the 4-bit and 8-bit implementations, BRAMs are preferentially consumed for weights and activations, and URAMs are utilized for the portions that exceed the BRAM capacity. Thus, BRAM utilizations of the 4-bit and 8-bit implementations are almost the same while their URAM utilizations differ. As a result, compared to the 8-bit quantization, the URAM utilization is further reduced by 17.7% with the aggressive 4-bit quantization. The number of DSP slices is the same in the 4-bit and 8-bit implementations regardless of the quantized bit widths. In this case, the benefit of the 4-bit implementation over the 8-bit one is the parameter size reduction that can save the scarce on-chip memory resources. Although our 8-bit FPGA implementation already meets the resource constraints of the Zynq UltraScale+ MPSoC ZCU104 platform, it would be possible to further reduce the number of DSP slices for the 4-bit implementation by adopting an aggressive DSP packing technique.
Figure 11: Best accuracy of each combination of quantized blocks (8-bit quant.)
Figure 12: Best accuracy of each combination of quantized blocks (4-bit quant.)
@@
Table 7: Top-1 accuracy in software and FPGA board
Platform Model 8-bit quant. (%) 4-bit quant. (%)
Simulation DS+Block3 81.23$\pm$0.27 80.03$\pm$0.13
FPGA board DS+Block3 79.68$\pm$0.37 76.97$\pm$0.36
@@ 6.2.5 FPGA inference accuracy
Here, our 4-bit and 8-bit FPGA implementations are executed on the ZCU104 platform and evaluated in terms of inference accuracy. As mentioned in Sec. 5.3, the DS+Block3 model is implemented on the ZCU104 platform. The 8-bit and 4-bit quantized models trained with learning rates of [5e-3, 1e-5] and [5e-4, 1e-5], respectively, are used for the implementation. Note that only DSBlocks and MHSABlock are quantized in the DS+Block3 model. The parameters for the other remaining blocks are represented as float32 in the software, and fixed-point in the FPGA implementation. To see this difference, the simulation and on-board results of the 4-bit and 8-bit implementations are compared in terms of inference accuracy. The mean accuracies in Table 7 demonstrate that the accuracy loss of the FPGA implementation is small in the 8-bit case. Note that if we can selectively implement the best parameters (among three training trials) in Table 5 on the FPGA, the maximum on-board accuracy in the 8-bit case is 80.20%; in this case, the on-board accuracy is comparable to that of the baseline floating-point model (i.e., 80.15%) which is denoted as “Pre-trained” in Table 5.